Solve the system using elimination. Make sure to show/explain your work for credit.
y = 2x - 9
x + 3y = 8

Answer by ankor@dixie-net.com(22740)

y = 2x - 9
x + 3y = 8
we want both equations in standard form change the 1st equation to
-2x + y = -9
x + 3y = 8
multiply the 2nd equation by 2
-2x + y = -9
2x + 6y = 16
------------------Addition eliminates x, find y
0 + 7y = 7
y = 7/7
y = 1
Find x using the 2nd original equation, replace y with 1
x + 3(1) = 8
x = 8 - 3
x = 5
:
Check solution of x=5, y=1 in the 1st original equation y = 2x - 9
1 = 2(5) - 9
1 = 10 - 9